OpenStudy (anonymous):

a parametric form of the equation of the line through the points A (1,1, -2) and B (-1,2,3) is:

OpenStudy (amistre64):

need a vector from A to B then

OpenStudy (amistre64):

given a vector <a,b,c> from A to B; attach it to any point in the line; they give 2 points to choose from x = xo + at y = yo + bt z = zo + zt